Supplemental Operations Report June 11th

Supplemental Operations Report 11 June 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine, but one day early/Martian Bio crew’s last day

Non-nominal systems: None

Action taken for non-nominal systems: N/A

Generator: Offline (for the most part.)

ScienceDome Dual Split: On AC at 74 degrees F during the heat of the day. Off at night. Will change to on 24/7 Sunday, June 12th

Solar— Operating 24/7

Solar— Nominal

Notes on power system: Nothing to report.

Propane Reading, station tank – not noted %

Propane Reading, director tank— not noted %

Propane Reading, intern tank— not noted %

Propane Reading, generator—not noted %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 300 gallons

Sojourner rover used: yes

Hours: 172.2

Beginning charge: 100

Ending charge: 80

Currently charging: yes

Curiosity rover is at 195.3 hours, charge is 93%.

Spirit rover is at 185.9, charge is 91%.

Percy is at 243.2 hours, charge is 82%.

Oppy overheated on the way to town, hours unknown but still under 100, charge is 94%.

Notes on rovers: All rovers have been moved to town for the summer.

ATV’s Used: (Honda, 350.1, 350.2, 300): Nothing to report

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: nothing to report

HabCar used and why, where? Yes, to town and Martian Bio study sites.

CrewCar used and why, where? Yes, to study sites for Martian Bio crew

Luna used and why, where? No. Luna is not on campus until fall.

Campus wide inspection, if action taken, what and why? Not done

General notes and comments: nothing to report

Summary of general operations: Winding down the season and getting things ready for the shutdown.

Summary of internet: Nothing to report

Summary of suits and radios: Radios and chargers have been stored on the lower deck in a cabinet. Batteries are in the lab so they will not overheat in the summer.

Summary of Hab operations: Crew food has been removed for the summer and all remaining food will be removed on Sunday.

Summary of Outpost operations: Intern trailer windows have been covered inside with reflective bubble wrap.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: AC on to 74 degrees F. Almost ready for shutdown.

Summary of RAM operations: Ready for shutdown

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: It’s very hot.

Copyright © The Mars Society. All rights reserved. | Main Site